# Enhancing Large Language Models’ Situated Faithfulness to External Contexts

### v1

@@ Abstract
Large Language Models (LLMs) are often augmented with external information as contexts, but this external information can sometimes be inaccurate or even intentionally misleading. We argue that robust LLMs should demonstrate situated faithfulness, dynamically calibrating their trust in external information based on their confidence in the internal knowledge and the external context. To benchmark this capability, we evaluate LLMs across several QA datasets, including a newly created dataset called RedditQA featuring in-the-wild incorrect contexts sourced from Reddit posts. We show that when provided with both correct and incorrect contexts, both open-source and proprietary models tend to overly rely on external information, regardless of its factual accuracy. To enhance situated faithfulness, we propose two approaches: Self-Guided Confidence Reasoning (SCR) and Rule-Based Confidence Reasoning (RCR). SCR enables models to self-access the confidence of external information relative to their own internal knowledge to produce the most accurate answer. RCR, in contrast, extracts explicit confidence signals from the LLM and determines the final answer using predefined rules. Our results show that for LLMs with strong reasoning capabilities, such as GPT-4o and GPT-4o mini, SCR outperforms RCR, achieving improvements of up to 24.2% over a direct input augmentation baseline. Conversely, for a smaller model like Llama-3-8B, RCR outperforms SCR. Fine-tuning SCR with our proposed Confidence Reasoning Direct Preference Optimization (CR-DPO) method improves performance on both seen and unseen datasets, yielding an average improvement of 8.9% on Llama-3-8B. In addition to quantitative results, we offer insights into the relative strengths of SCR and RCR. Our findings highlight promising avenues for improving situated faithfulness in LLMs. The data [^1] https://huggingface.co/datasets/kkkevinkkk/SituatedFaithfulnessEval and code [^2] https://github.com/kkkevinkkkkk/situated_faithfulness.git are released.
@@ 1 Introduction
Large Language Models (LLMs) can be enhanced by incorporating relevant external context (Chen et al., 2017; Lewis et al., 2020) provided by automated retrieval systems, tools, users, or another model (Mialon et al., 2023). However, external information can often be unreliable, containing errors due to online misinformation, intentional disinformation stemming from malice (Zou et al., 2024), noisy human inputs, and outdated data (Kasai et al., 2023). Prior research has demonstrated that LLMs can be too faithful to the external context, making them susceptible to being misled by unreliable contexts (Petroni et al., 2020; Xie et al., 2023; Shi et al., 2023; Shaier et al., 2024).
@@
Instead of blindly following external information, a robust LLM should exhibit situated faithfulness, dynamically balancing its reliance on internal knowledge and external context. This means the model should trust its internal knowledge when it is correct or when external information is unreliable, while conversely leaning on external data when uncertain or when the external evidence is convincing. The key objective is to maximize the accuracy of the model’s responses, irrespective of whether the external context is correct or incorrect. To achieve this, it requires strong reading comprehension skills to interpret and extract accurate information as well as confidence reasoning—the capacity to self-reflect on the confidence about the model’s own internal knowledge, and contrast it with the external context to arrive at the most accurate final answer.
@@
We benchmark the current LLM’s ability to perform situated faithful reasoning on several question-answering tasks, including FreshQA (Vu et al., 2023), NaturalQA (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023). These datasets reflect a diverse set of world knowledge across domains. To simulate both scenarios in which internal knowledge outperforms external evidence or vice versa, we pair each question with a correct context retrieved online and an incorrect context synthesized by LLMs. Besides leveraging automatically generated contexts of various kinds, we further examine how LLMs handle human-written incorrect contexts using RedditQA—a new dataset that we create by sourcing incorrect contexts from Reddit posts containing real-world factual mistakes that lead to incorrect answers to corresponding questions. Concurrent work, ClashEval (Wu et al., 2024), FaithEval (Ming et al., 2024) and DynamicQA (Marjanovi’c et al., 2024) benchmark a similar problem but don’t focus on solutions.
@@
We identify two classes of LLM-based approaches for confidence reasoning to improve situated faithfulness (Figure 1): self-guided confidence reasoning (SCR) and rule-based confidence reasoning (RCR). In SCR, the model is informed that external information might be incorrect and is allowed to reason either implicitly or explicitly via chain-of-thought to arrive at the final conclusion. In contrast, RCR first extracts model’s confidence signals regarding the internal knowledge and the external context, and then applies predefined rules to choose a source.
@@ 2 Related Work
Facing such knowledge conflicts(Xu et al., 2024), LLMs exhibit enigmatic behaviors. They can be too receptive to external evidence (Xie et al., 2024; Tan et al., 2024) and over-reliant on parametric knowledge (Longpre et al., 2021; Jin et al., 2024) at the same time. Shi et al. (2024) propose a fusion-in-decoder approach to trust context when sources are known to be reliable. Yu et al. (2024) resolve knowledge conflict by training a separate classifier with in-distribution data to remove untruthful information from context as a pre-processing step. However, it relies on in-distribution data and the model’s hidden states, limiting its practical use. Moreover, Doing so at the level of surface forms fails to consider various semantic granularities. Concurrent to our work, ClashEval (Wu et al., 2024), FaithEval (Ming et al., 2024) and DynamicQA (Marjanovi’c et al., 2024) also benchmark this problem but don’t focus on providing an effective solution.
It has been shown that LLMs are capable of estimating the confidence in the correctness of their own response (Kadavath et al., 2022; Xiong et al., 2023). In-context knowledge conflicts additionally require the model to corroborate the external with the internal to make reasoned comparisons to the final conclusion. Sharing perhaps the most similar idea to ours, concurrent work AstuteRAG (Wang et al., 2024) proposes to iteratively consolidate internal and external knowledge while considering source reliability. Their major setup leverages actual results from a commercial search engine, which presents an arbitrary mixture of correct, incorrect, and irrelevant evidence in a single context. Focusing on a clean evaluation of LLMs’ ability to handle a specific context, our setup controls the type of evidence the model sees at an instance level. We further explore various methods, including confidence-based, self-evaluation-based, and DPO methods, to elicit models’ ability to resolve knowledge conflicts–all of which perform effectively. Another concurrent work, KnowPO (Zhang et al., 2024), also employs DPO to teach models how to resolve knowledge conflicts. However, their approach uses preference pairs based solely on final answers, while our CR-DPO method extends this by incorporating COT reasoning before the final answer. This not only enhances the model’s reasoning ability but also provides explainability regarding its final decision.
@@ 3.1 Formulation
  • $\Pr(y(a)=1\mid y(c,q)=0,y(a_{in})=0)$: The probability of providing a correct answer when both the context and the internal answer are incorrect, which is unlikely to improve.
@@ 3.2.1 RedditQA
While many QA datasets are accompanied with correct human-generated documents containing valuable supporting evidence, no existing dataset focuses on human-written documents with inaccuracies, raising questions on how models react to natural false contexts. To fill in this gap, we contribute a new dataset sourced from existing Reddit posts, which often contain unverified, inaccurate claims. Concretely, we first apply a claim detector to Reddit post summaries ((Völske et al., 2017)), filtering out those without significant factual claims. Then, GPT-4o evaluates whether the claims may contain inaccuracies. For claims flagged as incorrect or uncertain, GPT-4o generates a self-contained, multiple-choice world-knowledge question, providing the correct answer, a wrong answer, and two plausible alternatives. Next, we use a natural language inference model to check if the incorrect post indeed provides the wrong answer to the question.
Following the automated process, we conduct a human evaluation to ensure the validity of each question by verifying the following criteria: 1) Is the question concise, self-contained, and asking for verifiable factual information based on world knowledge? 2) Is the answer confirmed by reputable sources, such as authoritative websites (e.g., government websites, Wikipedia) or textbooks? 3) Does the Reddit post contain inaccuracies that lead to an incorrect answer? 4) Are the other plausible answer choices clearly wrong and potentially misleading? 5) Is verified evidence from online sources correctly paired with the context for the QA instance? If any of these criteria are not met, annotators attempt to correct the errors. For non-correctable data points, annotators will discard them. Step by Step guidance and other annotation details can be found in §G.
@@ 3.2.2 Other QAs
  • ClashEval (Wu et al., 2024): a contemporary dataset with world knowledge questions from multiple domains, each paired with both correct and perturbed incorrect contexts. This dataset was available at the time of writing this paper, and we have included it as part of our evaluation.
For TriviaQA, NaturalQA, and FreshQA, we retrieve correct contexts from relevant websites and verify them with a natural language inference model. If needed, GPT-4o generates supporting contexts. Wrong contexts are created by having GPT-4o modify correct contexts to lead to incorrect answers. For PopQA, we use ConflictQA contexts with preprocessing to filter noise. Further details on evaluation and dataset statistics are in §A.
@@ 4.1 Self-Guided Confidence Reasoning
Implicit Self-Guided Confidence Reasoning (ImplicitSCR) The model is prompted that the provided context may be incorrect and must rely on its own judgment to assess the reliability of the context before providing a final answer. To encourage the use of internal knowledge, the context is presented after the question. Our pilot experiments (§E) indicate that this ordering structure biases the model to rely more on prior knowledge, reducing susceptibility to misleading context. After that, the model implicitly reasons about confidence during the decision-making process and outputs only the final answer. (See §F.4 for the actual prompts used).
Explicit Self-Guided Confidence Reasoning-Explicit (ExplicitSCR) the model first generates separate answers based on its internal knowledge (§F.1) and completely faithful to the provided context. (§F.3). The model then performs verbalized confidence reasoning using a chain-of-thought process. It begins by evaluating the confidence in its internal answer, reflecting on how it derived this answer based on known facts. Next, it assesses the reliability of the external context by cross-referencing it with the facts supporting its internal answer. The final answer is chosen by reasoning through both perspectives, balancing internal and contextual confidence. (See §F.5 for the details of the prompt.)
Confidence Reasoning Preference Optimization (CR-DPO) We train the model to learn verbalized confidence reasoning by optimizing its preferences between a pair of self-sampled reasoning paths. First, when the model’s internal answer is incorrect, we provide it with a correct external context and ask it to reason why the context is correct and its internal answer is wrong, resulting in a chosen reasoning path. Conversely, we lie to the model by telling it the context is incorrect and its internal answer is right, asking it to reason accordingly, generating a rejected reasoning path. When the model’s internal answer is correct, we repeat the same process, comparing correct internal reasoning with external context. (See §F.9 Prompts for these two processes)This approach allows the model to learn preferences between reasoning paths through direct preference optimization (Rafailov et al., 2023). Additionally, to encourage diversity in reasoning, we use dual sampling: for each instance, we sample two pairs of reasoning paths using different prompts and in-context examples. Following Pang et al. (2024), we apply negative log-likelihood to DPO loss to improve the optimization of the reasoning process. (See §B for implementation details)
@@ 4.2 Rule-based Confidence Reasoning
In this set of approaches, we define different types of confidence estimations for both LLMs’ internal and context-based answers. The confidence estimation can be represented by a numerical score or a self-evaluation result. The reasoning process is then carried out by predefined heuristics, which determine the final answer by comparing these confidence estimates. All methods here follow the same procedure of 1) generate an internal answer and a context answer (completely faithful to the contexts even if it’s wrong) separately; 2) they estimate confidence for each answer; 3) and utilize rules to select.
Internal Evaluation (InternalEval) The LLM self-evaluates the correctness of its internal answer with the prompt in §F.6. If the self-evaluation suggests the answer is correct, the internal answer is selected; otherwise, the context-based answer is used.
Context Evaluation (ContextEval) The model assesses the correctness of the context relative to the question (See §F.7 for the prompt). If the context is deemed correct, the context-based answer is used; otherwise, the internal answer is selected.
Internal Confidence Thresholding (InternalConf) The LLM evaluates its internal confidence in the internal answer. The reasoning rule here is that: if this confidence exceeds a predefined threshold, the internal answer is selected; otherwise, the model defaults to the context-based answer. The threshold could be prefined as 0.5, or it could be further tuned on a calibration set from the dev split of the same dataset if available. The confidence score could be sequence probability or other confidence elicitation methods like self-consistency.
@@ 4.3 Other Baselines
Direct Input Augmentation The simplest and yet common implementation of RAG systems is directly concatenating a question with the corresponding contexts. We follow the same structure of the generative prompt described in Yu et al. (2024) where a simple instruction is inputted first, following the external context concatenated with the question, forming a prompt that is then provided to the model. The model is instructed to leverage the external context in generating the answer. We utilize 3-shot to make sure the answer is in the right format. (See §F.2 for the prompt)
Truth-aware Context Selection Similar to context evaluation, Truth-aware context selection (Yu et al., 2024) filters out incorrect parts of the context at a granular sentence or token level using a classifier. Unlike rule-based methods that choose between internal and context-based answers, this approach feeds the filtered context back into the LLM to generate the final answer. However, it relies on hidden states, which are not accessible in proprietary models like GPT-4, and requires in-distribution training data, making it less comparable to methods that do not. Therefore, we adopt an alternative: LMs remove the untruthful sentences (TACS-LR) (See §F.8 for the prompt), and the refined context is used to produce the final answer.
@@ 5.2 LLMs can Do Both Self-Guided and Rule-Based Confidence Reasoning
Enhancing SCR Through CR-DPO: While SCR generally outperforms RCR methods on the GPT-4 series, it underperforms “InternalConf” on Llama-3-8B in Table 2. The most significant declines are in Accf, particularly in TriviaQA (64.6% → 34.0%) and RedditQA (75% → 43.8%). This indicates that while Llama-3-8B demonstrates some internal confidence estimation, it tends to favor external context when it is available. We attribute this bias in part to the model’s weaker reasoning abilities. As shown in Figure 2, when presented with misleading context, the ExplicitSCR model correctly lists the specific years Richard Daley served in Chicago. However, it acknowledges that it lacks information about him being mayor in another city. If the model possessed strong reasoning abilities, it should recognize that, regardless of whether Daley was mayor in another city, his 21-year tenure as mayor of Chicago makes Chicago the correct answer. By applying our proposed Confidence Reasoning Preference Optimization (CR-DPO) method, the model is able to detect the deceptive context and correctly identify the answer. CR-DPO significantly improves the model’s SCR performance, boosting it by +8.9% compared to ExplicitSCR(Table 2), and demonstrates robustness across various context types (§H).

### v2

# To Trust or Not to Trust? Enhancing Large Language Models’ Situated Faithfulness to External Contexts

@@ Abstract
Large Language Models (LLMs) are often augmented with external contexts, such as those used in retrieval-augmented generation (RAG). However, these contexts can be inaccurate or intentionally misleading, leading to conflicts with the model’s internal knowledge. We argue that robust LLMs should demonstrate situated faithfulness, dynamically calibrating their trust in external information based on their confidence in the internal knowledge and the external context to resolve knowledge conflicts. To benchmark this capability, we evaluate LLMs across several QA datasets, including a newly created dataset featuring in-the-wild incorrect contexts sourced from Reddit posts. We show that when provided with both correct and incorrect contexts, both open-source and proprietary models tend to overly rely on external information, regardless of its factual accuracy. To enhance situated faithfulness, we propose two approaches: Self-Guided Confidence Reasoning (SCR) and Rule-Based Confidence Reasoning (RCR). SCR enables models to self-assess the confidence of external information relative to their own internal knowledge to produce the most accurate answer. RCR, in contrast, extracts explicit confidence signals from the LLM and determines the final answer using predefined rules. Our results show that for LLMs with strong reasoning capabilities, such as GPT-4o and GPT-4o mini, SCR outperforms RCR, achieving improvements of up to 24.2% over a direct input augmentation baseline. Conversely, for a smaller model like Llama-3-8B, RCR outperforms SCR. Fine-tuning SCR with our proposed Confidence Reasoning Direct Preference Optimization (CR-DPO) method improves performance on both seen and unseen datasets, yielding an average improvement of 8.9% on Llama-3-8B. In addition to quantitative results, we offer insights into the relative strengths of SCR and RCR. Our findings highlight promising avenues for improving situated faithfulness in LLMs.
@@ 1 Introduction
Large Language Models (LLMs) can be enhanced by incorporating relevant external context (Chen et al., 2017; Lewis et al., 2020) provided by automated retrieval systems, tools, users, or another model (Mialon et al., 2023). However, external information can often be unreliable, containing errors due to online misinformation, intentional disinformation stemming from malice (Zou et al., 2024), noisy human inputs, and outdated data (Kasai et al., 2023), leading to knowledge conflicts with LLMs’ internal beliefs. Prior research has demonstrated that LLMs can be too faithful to the external context, making them susceptible to being misled by unreliable contexts (Petroni et al., 2020; Xie et al., 2023; Shi et al., 2023; Shaier et al., 2024).
@@
Instead of blindly following external information, a robust LLM should exhibit situated faithfulness, dynamically balancing its reliance on internal knowledge and external context to resolve knowledge conflict. This means the model should trust its internal knowledge when it is correct or when external information is unreliable, while conversely leaning on external data when uncertain or when the external evidence is convincing. The key objective is to maximize the accuracy of the model’s responses, irrespective of whether the external context is correct or incorrect. To achieve this, it requires strong reading comprehension skills to interpret and extract accurate information as well as confidence reasoning—the capacity to self-reflect on the confidence about the model’s own internal knowledge, and contrast it with the external context to arrive at the most accurate final answer.
@@
We benchmark the current LLM’s ability to perform situated faithful reasoning on several question-answering tasks, including FreshQA (Vu et al., 2023), NaturalQA (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023). These datasets reflect a diverse set of world knowledge across domains. To simulate both scenarios in which internal knowledge outperforms external evidence or vice versa, we pair each question with a correct context retrieved online and an incorrect context synthesized by LLMs. Besides leveraging automatically generated contexts of various kinds, we further examine how LLMs handle human-written incorrect contexts using RedditQA—a new dataset that we create by sourcing incorrect contexts from Reddit posts containing real-world factual mistakes that lead to incorrect answers to corresponding questions. Concurrent work, ClashEval (Wu et al., 2024), addresses a similar problem but uses only synthetically modified incorrect contexts and does not focus on an improvement methodology.
@@
We identify two classes of LLM-based approaches for confidence reasoning to improve situated faithfulness: self-guided confidence reasoning (SCR) and rule-based confidence reasoning (RCR). In SCR, the model is informed that external information might be incorrect and is allowed to reason either implicitly or explicitly via chain-of-thought to arrive at the final conclusion. In contrast, RCR first extracts model’s confidence signals regarding the internal knowledge and the external context, and then applies predefined rules to choose a source.
@@ 2 Related Work
Facing such knowledge conflicts, LLMs exhibit enigmatic behaviors. They can be too receptive to external evidence (Xie et al., 2024) and over-reliant on parametric knowledge (Longpre et al., 2021) at the same time. Shi et al. (2024) propose a fusion-in-decoder approach to trust context when sources are known to be reliable. Yu et al. (2024) resolve knowledge conflict by training a separate classifier to remove untruthful information from context as a pre-processing step. Doing so at the level of surface forms fail to consider various semantic granularities. Concurrent to our work, ClashEval (Wu et al., 2024), FaithEval (Ming et al., 2024) and DynamicQA (Marjanovi’c et al., 2024) also benchmark this problem but don’t focus on providing an effective solution.
LLMs can estimate confidence in their own responses (Kadavath et al., 2022; Xiong et al., 2023). Resolving in-context knowledge conflicts introduces additional challenges as it involves not only confidence estimation, but corroborating the external with internal to make reasoned comparison to the final conclusion. See more related work in Appendix F.
@@ 3.2.1 RedditQA
While many QA datasets are accompanied with correct human-generated documents containing valuable supporting evidence, no existing dataset focuses on human-written documents with inaccuracies, raising questions on how models react to natural false contexts. To fill in this gap, we contribute a new dataset sourced from existing Reddit posts, which often contain unverified, inaccurate claims. Concretely, we first apply a claim detector [^2] https://huggingface.co/whispAI/ClaimBuster-DeBERTaV2 to Reddit post summaries ((Völske et al., 2017)), filtering out those without significant factual claims. Then, GPT-4o evaluates whether the claims may contain inaccuracies. For claims flagged as incorrect or uncertain, GPT-4o generates a self-contained, multiple-choice world-knowledge question, providing the correct answer, a wrong answer, and two plausible alternatives. Next, we use a natural language inference model to check if the incorrect post indeed provides the wrong answer to the question.
Following the automated process, we conduct a human evaluation to ensure question validity by verifying: 1) Is the question concise, self-contained, and asking for verifiable factual information based on world knowledge? 2) Is the answer supported by authoritative sources (e.g., government websites, Wikipedia, textbooks)? 3) Does the Reddit post introduce inaccuracies leading to a wrong answer? 4) Are alternative answer choices clearly incorrect and misleading? 5) Is verified evidence from online sources correctly paired with the context for the QA instance? If any criteria are unmet, annotators attempt to correct errors. For non-correctable data points, annotators discard them. Step-by-step guidance and other annotation details can be found in §I.
@@ 3.2.2 Other QAs
For TriviaQA, NaturalQA, and FreshQA, we retrieve correct contexts from relevant websites and verify them with a natural language inference model. If needed, GPT-4o generates supporting contexts. Wrong contexts are created by having GPT-4o modify correct contexts to lead to incorrect answers. For PopQA, we use ConflictQA contexts with noise filtering. See §A for details.
@@ 4.1 Self-Guided Confidence Reasoning
Implicit Self-Guided Confidence Reasoning (ImplicitSCR) The model is prompted that the provided context may be incorrect and must rely on its own judgment to assess the reliability of the context before providing a final answer. To encourage the use of internal knowledge, the context is presented after the question. Our pilot experiments (§E) indicate that this ordering structure biases the model to rely more on prior knowledge, reducing susceptibility to misleading context. After that, the model implicitly reasons about confidence during the decision-making process and outputs only the final answer. (See §H.2 for the actual prompts used).
Explicit Self-Guided Confidence Reasoning-Explicit (ExplicitSCR) In this method, the model first generates separate answers based on its internal knowledge and the provided context (via two different prompts). The model then performs verbalized confidence reasoning using a chain-of-thought process. It begins by evaluating the confidence in its internal answer, reflecting on how it derived this answer based on known facts. Next, it assesses the reliability of the external context by cross-referencing it with the facts supporting its internal answer. The final answer is chosen by reasoning through both perspectives, balancing internal and contextual confidence. (See §H.3 for the details of the prompt.)
Confidence Reasoning Preference Optimization (CR-DPO) We train the model to learn verbalized confidence reasoning by optimizing its preferences between a pair of self-sampled reasoning paths. First, when the model’s internal answer is incorrect, we provide it with a correct external context and ask it to reason why the context is correct and its internal answer is wrong, resulting in a chosen reasoning path. Conversely, we lie to the model by telling it the context is incorrect and its internal answer is right, asking it to reason accordingly, generating a rejected reasoning path. When the model’s internal answer is correct, we repeat the same process, comparing correct internal reasoning with external context. (See §H.7 Prompts for these two processes)This approach allows the model to learn preferences between reasoning paths through direct preference optimization (Rafailov et al., 2023). Additionally, to encourage diversity in reasoning, we use dual sampling: for each instance, we sample two pairs of reasoning paths using different prompts and in-context examples. Following Pang et al. (2024), we apply negative log-likelihood to DPO loss to improve the optimization of the reasoning process. (See §B for implementation details)
@@ 4.2 Rule-based Confidence Reasoning
These approaches define confidence estimation for both internal and context-based answers, represented as numerical scores or self-evaluations. The reasoning process is then carried out by predefined heuristics, which determine the final answer by comparing these confidence estimates. All methods here follow the same procedure of 1) generate an internal answer and a context answer (completely faithful to the contexts even if it’s wrong) separately; 2) they estimate confidence for each answer; 3) and utilize rules to select.
Internal Evaluation (InternalEval) The LLM self-evaluates the correctness of its internal answer with the prompt in §H.4. If the self-evaluation suggests the answer is correct, the internal answer is selected; otherwise, the context-based answer is used.
Context Evaluation (ContextEval) The model assesses the correctness of the context relative to the question (See §H.5 for the prompt). If the context is deemed correct, the context-based answer is used; otherwise, the internal answer is selected.
Internal Confidence Thresholding (InternalConf) The LLM selects its internal answer if its confidence exceeds a predefined threshold; otherwise, it defaults to the context-based answer. The threshold can be set to 0.5 or tuned on a calibration set if available. Confidence can be derived from sequence probability or other elicitation methods like self-consistency. ActiveRAG (Jiang et al., 2023) reduces to InternalConf in non-iterative short-form generation. Unlike ActiveRAG’s token-level thresholds, our implementation uses answer-level thresholds, offering slight empirical advantages. See §D for a comparison.
@@ 4.3 Other Baselines
Direct Input Augmentation The simplest and yet common implementation of RAG systems is directly concatenating a question with the corresponding contexts. We follow the same structure of the generative prompt described in Yu et al. (2024) where a simple instruction is inputted first, following the external context concatenated with the question, forming a prompt that is then provided to the model. The model is instructed to leverage the external context in generating the answer. We utilize 3-shot to make sure the answer is in the right format. (See §H.1 for the prompt)
Truth Aware context-selection Similar to context evaluation, Truth-Aware context selection (Yu et al., 2024) filters out incorrect parts of the context at a granular sentence or token level using a classifier. Unlike rule-based methods that choose between internal and context-based answers, this approach feeds the filtered context back into the LLM to generate the final answer. However, it relies on hidden states, which are not accessible in proprietary models like GPT-4, and requires in-distribution training data, making it less comparable to methods that do not. Therefore, we adopt an alternative: LMs remove the untruthful sentences (TACS-LR) (See §H.6 for the prompt), and the refined context is used to produce the final answer.
@@ 5.2 LLMs can Do Both Self-Guided and Rule-Based Confidence Reasoning
Enhancing SCR Through CR-DPO: While SCR generally outperforms RCR methods on GPT-4 series, it underperforms “InternalConf” on Llama-3-8B (Table 2), with Accf dropping significantly in TriviaQA (64.6% → 34.0%) and RedditQA (75% → 43.8%). This indicates that while Llama-3-8B demonstrates some internal confidence estimation, it tends to favor external context when it is available. We attribute this bias in part to the model’s weaker reasoning abilities. As shown in Figure 2, when presented with misleading context, the ExplicitSCR model correctly lists the specific years Richard Daley served in Chicago. However, it acknowledges that it lacks information about him being mayor in another city. If the model possessed strong reasoning abilities, it should recognize that, regardless of whether Daley was mayor in another city, his 21-year tenure as mayor of Chicago makes Chicago the correct answer. By applying our proposed Confidence Reasoning Preference Optimization (CR-DPO) method, the model is able to detect the deceptive context and correctly identify the answer. CR-DPO significantly improves the model’s SCR performance, boosting it by +8.9% (Table 2), and demonstrates robustness across various context types (§J).